OpenStudy (anonymous):

Is this correct? (6-z)z= 6z-z^2

OpenStudy (unklerhaukus):

\[(6-z)z=6\times z-z\times z=6z-z^2\large\color{red}\checkmark\]

OpenStudy (anonymous):

Thank you. So I had it correct?